# NoteChat: A Dataset of Synthetic Patient-Physician Conversations Conditioned on Clinical Notes

### v2

@@ Abstract
We introduce NoteChat, a novel cooperative multi-agent framework leveraging Large Language Models (LLMs) to generate patient-physician dialogues. NoteChat embodies the principle that an ensemble of role-specific LLMs, through structured role-play and strategic prompting, can perform their assigned roles more effectively. The synergy among these role-playing LLMs results in a cohesive and efficient dialogue generation. Evaluation on MTS-dialogue Abacha et al. (2023); Ben Abacha et al. (2023), a benchmark dataset for patient-physician dialogues-note pairs, shows that models trained with the augmented synthetic patient-physician dialogues by NoteChat 111 Our synthetic patient-physician dialogue data is publicly available at https://huggingface.co/datasets/akemiH/NoteChat. outperforms other state-of-the-art models for generating clinical notes. Our comprehensive automatic and human evaluation demonstrates that NoteChat substantially surpasses state-of-the-art models like ChatGPT and GPT-4 up to 22.78% by domain experts in generating superior synthetic patient-physician dialogues based on clinical notes  [^2] Our codes and prompts are publicly available at https://github.com/believewhat/Dr.NoteAid.. NoteChat has the potential to engage patients directly and help clinical documentation, a leading cause of physician burnout Budd (2023).
@@ 1 Introduction
Figure 1: An illustration of NoteChat. Apricot indicates that our pipeline can generate smooth patient-physician conversations. Blue shows the characteristics of information seeking, where physicians can actively ask questions to advance the conversation, thanks to 2⃝ Roleplay module. In addition, compared with the corresponding note content, the generated utterances are more colloquial, but the key medical concepts are highly overlapped, which reflects NoteChat’s control over factuality (mainly from 1⃝ Planning module). Lavender means that NoteChat can generate reasonable explanations for patients, and a lot of information in the chat is reasonable imagination instead of hallucination. The two modules of 2⃝ Roleplay and 3⃝ Polish can stimulate the imaginative potential of LLMs and reduce unreasonable hallucination through self-examination.
@@ 3.2 Extrinsic Evaluation
Medical Chat Assistant: We used the PMC-Patient synthetic dialogues generated by ChatGPT, GPT4, and NoteChat to fine-tune the LLaMA2-7B  [^10] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, where we only used physician utterances as the training labels. Then, we evaluated these fine-tuned LLaMA2 chatbots on the ground truth dialogues from MTS-Dialog. For evaluation, recent work shows a higher human evaluation correlation for GPT-4 eval than traditional metrics Liu et al. (2023b); Gao et al. (2023); Fu et al. (2023); Zheng et al. (2023), so we also used the GPT4 preference as measurements to evaluate chatbots’ response quality. Specifically, we instruct GPT4 to give preference ranking  [^11] Prompts can be found in Appendix 8. based on the conversation history and the real response. We follow Yao et al. (2023a) to report the Mean Reciprocal Rank (MRR) Radev et al. (2002) of each model’s final ranking in Figure 2. Generally, a higher MRR implies that evaluators have a better alignment with the evaluators’ preferences.
@@ 5 Related Work
Clinical note and conversations generation: A task closely related to our work, but with an inverse direction, is the automatic generation of clinical notes from patient-physician conversations Krishna et al. (2020); Song et al. (2020); Yim and Yetisgen-Yildiz (2021); Su et al. (2022); Yao et al. (2023a). Recently, the MEDIQA-Chat 2023  [^18] https://sites.google.com/view/mediqa2023 introduced tasks in both directions (Dialogue2Note Summarization and Note2Dialogue Generation). However, their dataset is either private or limited to less than 2k examples. One of the main themes of recent data-centric AI is the synthetic data to overcome privacy concerns Pereira et al. (2022); Shafquat et al. (2022); Mishra et al. (2023). To the best of our knowledge, we are the first to introduce a large-scale publicly available patient-physician conversation dataset in English, each accompanied by corresponding medical documents, with an average number of utterances exceeding 20 rounds. In addition, our extrinsic eval shows that the NoteChat can be used as auxiliary data for Conversation2Note or Note2Conversation tasks and can also be used as a synthetic medical dialogue dataset alone to engage patients directly and help clinical documentation Zhang et al. (2023); Li et al. (2023b); Wang et al. (2023); Liu et al. (2023a); Xiong et al. (2023); Zeng et al. (2020).
@@ 7 Limitations and Ethical Considerations
Due to cost and time constraints, we could not try out many possibilities and alternatives in this paper. First of all, the current amount of data for human evaluation is not particularly sufficient. We are conducting more human evaluations. Secondly, due to cost issues, we currently do not use GPT-4 extensively to try the NoteChat pipeline. When OpenAI updates the Stateful API  [^19] https://www.reuters.com/technology/openai-plans-major-updates-lure-developers-with-lower-costs-sources-2023-10-11/, we will use this version to generate NoteChat-GPT4. Third, we extracted relevant UMLS-CUI codes for our Planning module, aiming to guide subsequent conversations around these critical terms. Such a checklist can help our pipeline improve factuality Asai et al. (2023); Huang et al. (2023), and can be very flexibly combined with other tools to meet different purposes, like information retrieval Khattab et al. (2022), entity&relation extraction Cai et al. (2023), medical jargon extraction Kwon et al. (2022), causal inference Yuan et al. (2023), evidence and reasoning path retrieval Asai et al. (2019, 2021), and many other knowledge injection ideas Fei et al. (2021); Yao and Yu (2021); Yao et al. (2023b).
@@ A.4 Color for Polish Promopt
Section
Subsection
Definition
@@
Chief Complaint
Patient’s primary motivation for the visit and type of visit
Review of Systems
Patient’s report of system related health and symptoms
Past Medical History
Patient’s reported diagnoses/conditions (when and what, excluding laboratory and imaging results and surgeries)
Past Surgical History
Patient’s reported prior surgeries (what, when, where)
Family Medical History
Conditions affecting patient’s close genetic relatives
Social History
Patient’s alcohol, tobacco, and drug related behaviors
Medications
Patient’s list of medications (not prescribed during visit)
Allergies
Patient’s list of allergies (primarily medicinal)
Miscellaneous
Patient’s clinically relevant social and other circumstances
@@
Immunizations
Vaccination record (not frequently discussed)
Laboratory and Imaging Results
Clinician’s discussion of laboratory/imaging results
@@
Assessment
Synthesis of reason for visit and pertinent diagnosis
@@
Diagnostics & Appointments
Plan for future tests, appointments, or surgeries
Prescriptions & Therapeutics
Plan for medications and therapeutics

### v3

@@ Abstract
We introduce NoteChat, a novel cooperative multi-agent framework leveraging Large Language Models (LLMs) to generate patient-physician dialogues. NoteChat embodies the principle that an ensemble of role-specific LLMs, through structured role-play and strategic prompting, can perform their assigned roles more effectively. The synergy among these role-playing LLMs results in a cohesive and efficient dialogue generation. Evaluation on MTS-dialogue Abacha et al. (2023); Ben Abacha et al. (2023), a benchmark dataset for patient-physician dialogues-note pairs, shows that models trained with the augmented synthetic patient-physician dialogues by NoteChat 111 Our synthetic patient-physician dialogue data is in supplementary material and are publicly available together with all codes and prompts at github:https://github.com/believewhat/Dr.NoteAid. outperforms other state-of-the-art models for generating clinical notes. Our comprehensive automatic and human evaluation demonstrates that NoteChat substantially surpasses state-of-the-art models like ChatGPT and GPT-4 up to 22.78% by domain experts in generating superior synthetic patient-physician dialogues based on clinical notes. NoteChat has the potential to engage patients directly and help clinical documentation, a leading cause of physician burnout Budd (2023).
@@ 1 Introduction
Figure 1: An illustration of NoteChat. Apricot indicates that our pipeline can generate smooth patient-physician conversations. Blue shows the characteristics of information seeking, where physicians can actively ask questions to advance the conversation, thanks to ② Roleplay module. In addition, compared with the corresponding note content, the generated utterances are more colloquial, but the key medical concepts are highly overlapped, which reflects NoteChat’s control over factuality (mainly from ① Planning module). Lavender means that NoteChat can generate reasonable explanations for patients, and a lot of information in the chat is reasonable imagination instead of hallucination. The two modules of ② Roleplay and ③ Polish can stimulate the imaginative potential of LLMs and reduce unreasonable hallucination through self-examination.
@@ 3.2 Extrinsic Evaluation
Medical Chat Assistant: We used the PMC-Patient synthetic dialogues generated by ChatGPT, GPT4, and NoteChat to fine-tune the LLaMA2-7B  [^9] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, where we only used physician utterances as the training labels. Then, we evaluated these fine-tuned LLaMA2 chatbots on the ground truth dialogues from MTS-Dialog. For evaluation, recent work shows a higher human evaluation correlation for GPT-4 eval than traditional metrics Liu et al. (2023b); Gao et al. (2023); Fu et al. (2023); Zheng et al. (2023), so we also used the GPT4 preference as measurements to evaluate chatbots’ response quality. Specifically, we instruct GPT4 to give preference ranking  [^10] Prompts can be found in Appendix 8. based on the conversation history and the real response. We follow Yao et al. (2023) to report the Mean Reciprocal Rank (MRR) Radev et al. (2002) of each model’s final ranking in Figure 2. Generally, a higher MRR implies that evaluators have a better alignment with the evaluators’ preferences.
@@ 5 Related Work
Clinical note and conversations generation: A task closely related to our work, but with an inverse direction, is the automatic generation of clinical notes from patient-physician conversations Krishna et al. (2020); Song et al. (2020); Yim and Yetisgen-Yildiz (2021); Su et al. (2022); Yao et al. (2023). Recently, the MEDIQA-Chat 2023  [^17] https://sites.google.com/view/mediqa2023 introduced tasks in both directions (Dialogue2Note Summarization and Note2Dialogue Generation). However, their dataset is either private or limited to less than 2k examples. One of the main themes of recent data-centric AI is the synthetic data to overcome privacy concerns Pereira et al. (2022); Shafquat et al. (2022); Mishra et al. (2023). To the best of our knowledge, we are the first to introduce a large-scale publicly available patient-physician conversation dataset in English, each accompanied by corresponding medical documents, with an average number of utterances exceeding 20 rounds. In addition, our extrinsic eval shows that the NoteChat can be used as auxiliary data for Conversation2Note or Note2Conversation tasks and can also be used as a synthetic medical dialogue dataset alone to engage patients directly and help clinical documentation Zhang et al. (2023); Li et al. (2023b); Wang et al. (2023); Liu et al. (2023a); Xiong et al. (2023); Zeng et al. (2020).
@@ 7 Limitations and Ethical Considerations
Due to cost and time constraints, we could not try out many possibilities and alternatives in this paper. First of all, the current amount of data for human evaluation is not particularly sufficient. We are conducting more human evaluations. Secondly, due to cost issues, we currently do not use GPT-4 extensively to try the NoteChat pipeline. When OpenAI updates the Stateful API  [^18] https://www.reuters.com/technology/openai-plans-major-updates-lure-developers-with-lower-costs-sources-2023-10-11/, we will use this version to generate NoteChat-GPT4. Third, we extracted relevant UMLS-CUI codes for our Planning module, aiming to guide subsequent conversations around these critical terms. Such a checklist can help our pipeline improve factuality Asai et al. (2023); Huang et al. (2023), and can be very flexibly combined with other tools to meet different purposes, like information retrieval Khattab et al. (2022), entity&relation extraction Cai et al. (2023), medical jargon extraction Kwon et al. (2022), causal inference Yuan et al. (2023), evidence and reasoning path retrieval Asai et al. (2019, 2021), and many other knowledge injection ideas Fei et al. (2021); Yao and Yu (2021).
@@ A.4 Color for Polish Promopt
Section Subsection Definition
@@
Chief Complaint Patient’s primary motivation for the visit and type of visit
Review of Systems Patient’s report of system related health and symptoms
Past Medical History Patient’s reported diagnoses/conditions (when and what, excluding laboratory and imaging results and surgeries)
Past Surgical History Patient’s reported prior surgeries (what, when, where)
Family Medical History Conditions affecting patient’s close genetic relatives
Social History Patient’s alcohol, tobacco, and drug related behaviors
Medications Patient’s list of medications (not prescribed during visit)
Allergies Patient’s list of allergies (primarily medicinal)
Miscellaneous Patient’s clinically relevant social and other circumstances
@@
Immunizations Vaccination record (not frequently discussed)
Laboratory and Imaging Results Clinician’s discussion of laboratory/imaging results
@@
Assessment Synthesis of reason for visit and pertinent diagnosis
@@
Diagnostics & Appointments Plan for future tests, appointments, or surgeries
Prescriptions & Therapeutics Plan for medications and therapeutics
@@
A.5 Ablation Study for Planning Module
To demonstrate the importance of the planning module, we designed the following experiment:
We conducted evaluations using both GPT-4 and human assessments. In the absence of the checklist and planning module, relying solely on role play and polishing for dialogue generation, the results were as follows:
  • GPT-4 Evaluation Win Rate:
    • Our model (without checklist & planning): 32%
    • GPT-4: 68%
  • Human Evaluation Win Rate:
    • Our model (without checklist & planning): 38%
    • GPT-4: 62%
The absence of the checklist and planning module resulted in the model’s inability to ensure comprehensive coverage of necessary information. While the generated dialogues were logically coherent, they significantly lacked informational content. This deficiency is primarily attributable to our model being based on GPT-3.5, which has a substantially lower capacity for information coverage compared to GPT-4.
Furthermore, when relying solely on a randomly ordered checklist, the results were as follows:
  • GPT-4 Evaluation Win Rate:
    • Our model (without planning module): 54%
    • GPT-4: 46%
  • Human Evaluation Win Rate:
    • Our model (without planning module): 40%
    • GPT-4: 60%
These results indicate slight differences. When evaluated by GPT-4, our model without the planning module appeared superior due to providing more information in shorter dialogue turns and extended conversations. However, human evaluators found the generated dialogues logically disorganized, primarily due to the absence of the planning module. The randomly ordered checklist led to each conversational turn lacking logical progression, making it seem less like a real dialogue. This highlights the critical importance of the planning module.